undeadknight26 (undeadknight26):

The table below shows the distance d(t) in feet that an object travels in t seconds. t (second) d(t) (feet) 3 126 4 224 5 350 6 504 What is the average rate of change of d(t) between 3 seconds and 5 seconds and what does it represent?

undeadknight26 (undeadknight26):

59.5 m/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 m/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 m/s; it represents the average distance traveled by the object between 3 seconds and 5 seconds 59.5 m/s; it represents the average distance traveled by the object between 3 seconds and 5 seconds

OpenStudy (kewlgeek555):

Well...\[Average Rate of Change=\frac{ f(b)-f(a) }{ b-a }\]Where f(b) is the y-value for 3 seconds (126), f(a) is the x-value for 3 seconds (3), b is the y-value for 5 seconds (350), and a is the x-value for 5 seconds (5)/

OpenStudy (kewlgeek555):

NO! It's the other way around. Sorry! Hold on!

OpenStudy (kewlgeek555):

f(b) is 350 and f(a) is 5. b is 126 and a is 3.

OpenStudy (kewlgeek555):

Sorry for the confusion.

OpenStudy (kewlgeek555):

Anyways, this is your new equation:\[Average Rate of Change=\frac{ 350-5 }{ 126-3 }\]

OpenStudy (kewlgeek555):

Solve that and you've got your answer. :)

OpenStudy (kewlgeek555):

AND it represents the average speed of the object between 3 seconds and 5 seconds.

undeadknight26 (undeadknight26):

345/123?

OpenStudy (kewlgeek555):

Yep. :) Simplify now.

undeadknight26 (undeadknight26):

2.8!

OpenStudy (kewlgeek555):

Hm.

undeadknight26 (undeadknight26):

oh simplify i divided lol...

OpenStudy (kewlgeek555):

No, thats the same thing.

undeadknight26 (undeadknight26):

oh ok...

OpenStudy (kewlgeek555):

Wow. Hm....

OpenStudy (kewlgeek555):

OH!

OpenStudy (kewlgeek555):

I think I'VE GOT IT.

OpenStudy (kewlgeek555):

\[Average Rate of Change=\frac{ 350-126 }{ 5-3 }\]

undeadknight26 (undeadknight26):

ahhhh!!!

undeadknight26 (undeadknight26):

112!

OpenStudy (kewlgeek555):

Okay, there's your answer. Lol.

OpenStudy (kewlgeek555):

Sorry for the confusion.

undeadknight26 (undeadknight26):

np!

undeadknight26 (undeadknight26):

so wait is it b?

undeadknight26 (undeadknight26):

@kewlgeek555 please help...

OpenStudy (kewlgeek555):

Yes, I believe it's B.

undeadknight26 (undeadknight26):

k thanks! neeed any help with anything?